# IntelliRadar: A Comprehensive Platform to Pinpoint Malicious Packages Information from Cyber Intelligence

### v4

@@ 4.3. RQ1: Effectiveness
CTIKG (huang2024ctikg, ) 53.85% 12.44% 20.22% 591 72
SecBERT (host2023constructing, ) 0.10% 2.07% 0.20% 661 13,340
@@
To address potential data leakage concerns where the malicious intelligence might have been encountered during the training phase of these models, we conducted an additional evaluation on fresh data. We randomly selected 50 security intelligence webpages published after January 1, 2025, ensuring that all content postdates the training cutoff of the evaluated models. As shown in Table 5, IntelliRadar achieves the best overall performance with 93.70% F1-score, with only 1 false positive indicating strong precision maintenance on unseen data. These findings on unseen data align with our evaluation results (Table 4), demonstrating that IntelliRadar consistently outperforms baseline models regardless of whether the malicious package intelligence was encountered during training. In addition to our core task of extracting malicious package names and versions, we also evaluated our approach on extracting other supplementary security-related entities. As shown in Table 6, while our primary contribution targets package identification, our method also demonstrates good performance on additional entities such as IOC (87.10% F1) and Discoverer (87.32% F1) extraction.
@@ 4.5. RQ3: Timeliness
Compared to the Snyk database. For 10,314 NPM malicious packages, IntelliRadar recorded the intelligence earlier than Snyk, accounting for 76.6%; among these, 6,598 packages were recorded 100.3 days earlier, with 49% of packages recorded a year in advance. For PyPI malicious packages, 5,765 packages were recorded earlier than Snyk, accounting for 70.3%, with 5,196 packages recorded more than a week earlier. Additionally, 566 malicious packages had consistent recording times, and only 189 packages (2.3%) were recorded earlier by Snyk. The earlier detection capabilities of IntelliRadar stem from our comprehensive coverage of social media and unstructured websites from security companies, enabling us to discover the latest malicious packages immediately. For the packages that Snyk detected earlier, their internal proprietary detection tools identified these specific threats before public disclosure. These delays in Snyk are attributed to its dependency on established databases, GitHub security advisories, and manual research processes (snykvulndb, ).
@@
Compared to the OSV database. IntelliRadar and OSV have consistent recording times for 15,479 NPM packages, accounting for 86.0% of the total. Additionally, IntelliRadar recorded 171 malicious packages earlier than OSV, primarily collected from Sonatype and Securityaffairs (securityaffairs_blog_ref, ). For PyPI malicious packages, IntelliRadar recorded 4,711 earlier than OSV, with most differences within a week; 53.7% were just one day earlier, mainly from Phylum. 511 packages were recorded on the same day by both IntelliRadar and OSV. Furthermore, we extracted PyPI package release times from Google’s bigquery-public-data and NPM package release times from the NPM Registry. Analysis shows that 4,533 NPM and 549 PyPI malicious package names were identified by IntelliRadar on the day of their release, indicating our method can swiftly collect relevant intelligence once malicious packages are publicly reported. OSV’s delays are attributed to its aggregation model, which depends on multiple upstream databases (osvdata, ) and integrated detection tools (packageanalysis, ).
@@ 4.7. RQ5: Usability
Package Name Versions PyPI Mirrors Database Downloads Tsinghua Tencent Aliyun Douban BFSU OSV Snyk GitHub Advisory IntelliRadar (Our) 1inch 8.6 - 8.9 - - - 10/10/22 (+3) 07/10/22 110 libcontroltoolver 4.86 - - - 26/02/23 (+0) 07/03/23 (+9) 26/02/23 62 matplotlyib 1.0.0 - - 29/03/23 (+31) 26/02/23 64 pipcryptov4 1.0.0 - - - 25/06/24 (+266) 05/10/23 (+2) 03/10/23 301 libideeee 1.0.0 - - - 25/06/24 (+294) 07/09/23 (+2) 05/09/23 74 ethereum2 2.8.4, 2.8.6, 2.8.9 - - - 17/12/23 (+436) 07/10/22 331 gkjzjh146 1.3 - - - 14/05/23 (+226) 30/09/22 638 httprequesthub 2.31.0 - 2.31.4 - - - 25/06/24 (+186) 25/12/23 (+3) 22/12/23 496 logic2 0.1.4 - - - - 29/03/23 (+104) 15/12/22 520 pipsqlite3liberyV2 1.1.0 - - - - - 22/05/23 (+104) 15/05/23 145 flak7 4.5.2 - - - - 07/09/22 (+5) 02/09/22 18 simpeljson 4.5.2 - - - - 11/02/23 20/06/23 (+129) 11/02/23 67 pyward 3.0 - - - - 08/09/23 (+1) 07/09/23 559 studypong 5.66, 7.16, 8.22, 10.45 - - - - 25/02/23 (+1) 07/03/23 (+11) 24/02/23 264 reqkests 2.28.1 - - - - 11/12/22 (+2) 09/12/22 143 beautiflulsoup 1.0.0 - - - 29/03/24 (+1) 28/03/24 76 pycryptdome 4.4.2 - - - - - 25/08/22 63 1337z 4.4.7 - - - - 31/08/22 176 urllib7 1.26.12 - - - - 15/12/22 69 urllib12 1.26.12, 1.30.0 - - - - 15/12/22 296
@@
To this end, we scanned the main downstream PyPI mirrors (guo2023empirical, ) using the intelligence database. As of 01/05/2025, we discovered 1,981 malicious packages across various PyPI mirrors, specifically 781 in the Tsinghua mirror, 548 in the Tencent mirror, 395 in the Douban mirror, 21 in the BFSU mirror, 132 in the Huawei mirror, and 104 in the Aliyun mirror. Among these malicious packages, 54.7% were uniquely identified in our database. These malicious packages had been downloaded over 86,240 times cumulatively, with the most severe case being the ethereum2 package, which had been downloaded 436 times before our identification and was identified 436 days earlier than other security databases such as Snyk. Table  7 presents the most impactful cases, highlighting the advantage of our tool in early identification.
@@ 5. Discussion
$\bullet$ Intelligence Collection Beyond PyPI and NPM. In this study, we successfully collected a large amount of malicious package intelligence related to PyPI and NPM, as well as 928 other types of threat intelligence, covering component vulnerabilities, CVE vulnerabilities, and malicious package information from other package managers. It is worth noting that we discovered malicious components Prettier in the VSCode Marketplace (prettier_vscode, ), which uses typesquatting for attacks (prettier_vscode, ). We also identified the XZ Utils software supply chain attack case. Our approach can quickly adapt to the intelligence requirements of various package manager platforms and effectively expand to new security domains, providing the security community with a more timely automated intelligence discovery.
$\bullet$ Data Poisoning Attacks. To combat potential data poisoning attacks, our framework implements multiple defensive mechanisms across various stages of the intelligence processing pipeline. During the source identification phase, we meticulously screen and validate intelligence sources to ensure their strong correlation with package management ecosystems, significantly mitigating the risk of incorporating contaminated data from unreliable sources. The intelligence aggregation mechanism serves as an additional safeguard, employing a voting system to cross-validate entities extracted from multiple independent sources, effectively minimizing the impact of potentially poisoned data from any single source. The intelligence extraction process further strengthens our defense through a three-step CoT approach—comprising entity extraction, relationship alignment, and verification—which facilitates the identification and filtration of inconsistent or suspicious intelligence. The effectiveness of our approach is strongly validated by experimental results: among 34,313 pieces of intelligence collected regarding NPM and PyPI packages, 29,485 (85.9%) were corroborated against authoritative security databases including OSV, Snyk and GitHub Advisory, which undergo rigorous verification by security researchers. This high correspondence rate with authoritative sources demonstrates that our multilayered defense strategy successfully maintains the integrity and reliability of collected intelligence while effectively countering data poisoning attempts.
@@ 6.1. Open Source Intelligence Analysis
Open-source intelligence (OSINT) collects information through public channels, widely used in cybersecurity. Researchers have developed automated systems to gather threat intelligence from various sources, including the internet (mulwad2011extracting, ), social media (sabottke2015vulnerability, ), developer communities (purba2023extracting, ), security forums, and code repositories (neil2018mining, ; bouwman2022helping, ; huang2021hackerrank, ). Recent work has applied LLMs for security intelligence analysis, with CTIKG (huang2024ctikg, ) using LLMs for knowledge graph construction from cyber threat intelligence, and Høst et al. (hostconstructing, ) constructing knowledge graphs from vulnerability descriptions in the NVD. In software engineering, extensive research has focused on named entity recognition (NER) and relation extraction (RE) from unstructured text sources. Studies have extracted software entities from tweets (zhang2022benchmarking, ), Stack Overflow posts (huang2023api, ; huo2022arclin, ), and technical documents (nguyen2023software, ). These approaches have been applied to API mention resolution (huang2023api, ), crash solutions (zhang2022benchmarking, ), library recognition (zhang2022benchmarking, ), and version incompatibility detection (zhao2023knowledge, ). However, malicious package intelligence extraction presents unique challenges distinct from traditional software entity recognition. IntelliRadar requires: (1) distinguishing malicious package names from benign packages mentioned in the same text (where benign packages are often the legitimate targets that malicious packages attempt to mimic), (2) identifying the specific ecosystem (npm vs PyPI) for each package, and (3) performing semantic-level understanding to extract complex attack vectors and indicators of compromise.
@@ 6.2. Software Supply Chain Security
Software supply chain security faces severe challenges, with package management systems becoming hotspots for malicious activities. Attackers inject malicious code through methods like code tampering, dependency confusion, and typesquatting, leading to data leakage and system damage. DL Vu et al. reveal various attack methods and anti-detection techniques (vu2023bad, ; duan2020towards, ; gu2023investigating, ). To address these threats, researchers have employed strategies ranging from static and dynamic analysis to machine learning. The MalOSS framework proposed by Duan et al. (duan2020towards, ) demonstrates similarities in malicious behaviors across different languages through metadata and API call sequence analysis. Huang et al. utilize graph-based behavior modeling to detect malicious packages (huang2024spiderscan, ). Zhang et al.(zhang2023malicious, ) utilize deep learning to understand malicious software’s behavioral semantics, while Liang et al.(liang2023needle, ) demonstrate the effectiveness of anomaly analysis algorithms. Additionally, GitHub, as the largest source code repository, may serve as a channel for distributing malicious code, with inconsistencies between distributed packages and source code potentially indicating malicious injection (vu2021lastpymile, ). However, existing detection methods face a critical limitation: detection results are scattered across security blogs and unstructured webpages, failing to reach downstream users effectively. Even with numerous detection tools, malicious packages persist because intelligence remains fragmented. IntelliRadar addresses this through a fundamentally different approach—systematically analyzing malicious package intelligence from dispersed sources, covering both historical and newly reported packages.
@@ 7. Conclusion
In the open-source ecosystem, package managers like PyPI and NPM lack up-to-date intelligence databases, allowing malicious packages to persist. To address this issue, we developed IntelliRadar, a system that leverages multi-source data collection and LLMs to gather 34,313 pieces of intelligence about NPM and PyPI from 24 sources, offering unique and detailed insights. The system enables earlier detection of malicious packages, reducing their impact, with a cost-effective rate of \$0.003 per intelligence item. This work presents an efficient and economical solution to enhance open-source software supply chain security.

### v6

@@ 4.3. RQ1: Effectiveness
CTIKG (huang2024ctikg) 53.85% 12.44% 20.22% 591 72
SecBERT (host2023constructing) 0.10% 2.07% 0.20% 661 13,340
@@
To address data leakage concerns, we conducted an additional evaluation on 50 security intelligence webpages published after January 1, 2025, ensuring all content postdates the training cutoff of evaluated models. As shown in Table 4, IntelliRadar achieves 93.70% F1-score with only 1 false positive, indicating strong precision on unseen data. These findings align with our evaluation results (Table 3), demonstrating that IntelliRadar consistently outperforms baseline models on both seen and unseen data. Beyond our core task of extracting package names and versions, we also evaluated other security-related entities. As shown in Table 5, our method demonstrates good performance on supplementary entities such as IOC (87.10% F1) and Discoverer (87.32% F1) extraction.
@@ 4.5. RQ3: Timeliness
Compared to the Snyk database. For 10,314 NPM malicious packages, IntelliRadar recorded the intelligence earlier than Snyk, accounting for 76.6%; among these, 6,598 packages were recorded 100.3 days earlier, with 49% of packages recorded a year in advance. For PyPI malicious packages, 5,765 packages were recorded earlier than Snyk, accounting for 70.3%, with 5,196 packages recorded more than a week earlier. Additionally, 566 malicious packages had consistent recording times, and only 189 packages (2.3%) were recorded earlier by Snyk. The earlier detection capabilities of IntelliRadar stem from our comprehensive coverage of social media and unstructured websites from security companies, enabling us to discover the latest malicious packages immediately. For the packages that Snyk detected earlier, their internal proprietary detection tools identified these specific threats before public disclosure. These delays in Snyk are attributed to its dependency on established databases, GitHub security advisories, and manual research processes (snykvulndb).
@@
Compared to the OSV database. IntelliRadar and OSV have consistent recording times for 15,479 NPM packages, accounting for 86.0% of the total. Additionally, IntelliRadar recorded 171 malicious packages earlier than OSV, primarily collected from Sonatype and Securityaffairs (securityaffairs_blog_ref). For PyPI malicious packages, IntelliRadar recorded 4,711 earlier than OSV, with most differences within a week; 53.7% were just one day earlier, mainly from Phylum. 511 packages were recorded on the same day by both IntelliRadar and OSV. Furthermore, we extracted PyPI package release times from Google’s bigquery-public-data and NPM package release times from the NPM Registry. Analysis shows that 4,533 NPM and 549 PyPI malicious package names were identified by IntelliRadar on the day of their release, indicating our method can swiftly collect relevant intelligence once malicious packages are publicly reported. OSV’s delays are attributed to its aggregation model, which depends on multiple upstream databases (osvdata) and integrated detection tools (packageanalysis).
@@ 4.7. RQ5: Usability
Package Name Versions PyPI Mirrors Database Downloads
Tsinghua Tencent Aliyun Douban BFSU OSV Snyk GitHub Advisory IntelliRadar (Our)
1inch 8.6 - 8.9 - - - 10/10/22 (+3) 07/10/22 110
libcontroltoolver 4.86 - - - 26/02/23 (+0) 07/03/23 (+9) 26/02/23 62
matplotlyib 1.0.0 - - 29/03/23 (+31) 26/02/23 64
pipcryptov4 1.0.0 - - - 25/06/24 (+266) 05/10/23 (+2) 03/10/23 301
libideeee 1.0.0 - - - 25/06/24 (+294) 07/09/23 (+2) 05/09/23 74
ethereum2 2.8.4, 2.8.6, 2.8.9 - - - 17/12/23 (+436) 07/10/22 331
gkjzjh146 1.3 - - - 14/05/23 (+226) 30/09/22 638
httprequesthub 2.31.0 - 2.31.4 - - - 25/06/24 (+186) 25/12/23 (+3) 22/12/23 496
logic2 0.1.4 - - - - 29/03/23 (+104) 15/12/22 520
pipsqlite3liberyV2 1.1.0 - - - - - 22/05/23 (+104) 15/05/23 145
flak7 4.5.2 - - - - 07/09/22 (+5) 02/09/22 18
simpeljson 4.5.2 - - - - 11/02/23 20/06/23 (+129) 11/02/23 67
pyward 3.0 - - - - 08/09/23 (+1) 07/09/23 559
studypong 5.66, 7.16, 8.22, 10.45 - - - - 25/02/23 (+1) 07/03/23 (+11) 24/02/23 264
reqkests 2.28.1 - - - - 11/12/22 (+2) 09/12/22 143
beautiflulsoup 1.0.0 - - - 29/03/24 (+1) 28/03/24 76
pycryptdome 4.4.2 - - - - - 25/08/22 63
1337z 4.4.7 - - - - 31/08/22 176
urllib7 1.26.12 - - - - 15/12/22 69
urllib12 1.26.12, 1.30.0 - - - - 15/12/22 296
@@
To this end, we scanned the main downstream PyPI mirrors (guo2023empirical) using the intelligence database. As of 01/05/2025, we discovered 1,981 malicious packages across various PyPI mirrors, specifically 781 in the Tsinghua mirror, 548 in the Tencent mirror, 395 in the Douban mirror, 21 in the BFSU mirror, 132 in the Huawei mirror, and 104 in the Aliyun mirror. Among these malicious packages, 54.7% were uniquely identified in our database. These malicious packages had been downloaded over 86,240 times cumulatively, with the most severe case being the ethereum2 package, which had been downloaded 436 times before our identification and was identified 436 days earlier than other security databases such as Snyk. Table  6 presents the most impactful cases, highlighting the advantage of our tool in early identification.
@@ 5. Discussion
$\bullet$ Intelligence Collection Beyond PyPI and NPM. In this study, we successfully collected a large amount of malicious package intelligence related to PyPI and NPM, as well as 928 other types of threat intelligence, covering component vulnerabilities, CVE vulnerabilities, and malicious package information from other package managers. It is worth noting that we discovered malicious components Prettier in the VSCode Marketplace (prettier_vscode), which uses typesquatting for attacks (prettier_vscode). We also identified the XZ Utils software supply chain attack case. Our approach can quickly adapt to the intelligence requirements of various package manager platforms and effectively expand to new security domains, providing the security community with a more timely automated intelligence discovery.
$\bullet$ Data Poisoning Attacks. To combat potential data poisoning attacks, our framework implements multiple defensive mechanisms across the intelligence processing pipeline. During source identification, we meticulously screen and validate intelligence sources to ensure strong correlation with package management ecosystems, mitigating the risk of contaminated data. The intelligence aggregation mechanism employs a voting system to cross-validate entities from multiple independent sources, minimizing the impact of poisoned data. The intelligence extraction process strengthens defense through a three-step CoT approach (entity extraction, relationship alignment, and verification) that identifies and filters inconsistent or suspicious intelligence. The effectiveness is validated by experimental results: among 34,313 pieces of intelligence collected, 29,485 (85.9%) were corroborated against authoritative security databases (OSV, Snyk, GitHub Advisory). This high correspondence rate demonstrates that our multilayered defense strategy successfully maintains the integrity and reliability of collected intelligence while effectively countering data poisoning attempts.
@@ 6.1. Open Source Intelligence Analysis
Open-source intelligence (OSINT) collects information through public channels, widely used in cybersecurity. Researchers have developed automated systems to gather threat intelligence from various sources, including the internet (mulwad2011extracting), social media (sabottke2015vulnerability), developer communities (purba2023extracting), security forums, and code repositories (neil2018mining; bouwman2022helping; huang2021hackerrank). Recent work has applied LLMs for security intelligence analysis (guo2025automated), with CTIKG (huang2024ctikg) using LLMs for knowledge graph construction from cyber threat intelligence, and Høst et al. (hostconstructing) constructing knowledge graphs from vulnerability descriptions in the NVD. In software engineering, extensive research has focused on named entity recognition (NER) and relation extraction (RE) from unstructured text sources. Studies have extracted software entities from tweets (zhang2022benchmarking), Stack Overflow posts (huang2023api; huo2022arclin), and technical documents (nguyen2023software), with recent work exploring self-evolving software intelligence management (liu2025evolaris). These approaches have been applied to API mention resolution (huang2023api), crash solutions (zhang2022benchmarking), library recognition (zhang2022benchmarking), and version incompatibility detection (zhao2023knowledge). However, malicious package intelligence extraction presents unique challenges distinct from traditional software entity recognition. IntelliRadar requires: (1) distinguishing malicious package names from benign packages mentioned in the same text (where benign packages are often the legitimate targets that malicious packages attempt to mimic), (2) identifying the specific ecosystem (npm vs PyPI) for each package, and (3) performing semantic-level understanding to extract complex attack vectors and indicators of compromise.
@@ 6.2. Software Supply Chain Security
Software supply chain security faces severe challenges, with package management systems becoming hotspots for malicious activities. Attackers inject malicious code through methods like code tampering, dependency confusion, and typesquatting, leading to data leakage and system damage. DL Vu et al. reveal various attack methods and anti-detection techniques (vu2023bad; duan2020towards; gu2023investigating). To address these threats, researchers have employed strategies ranging from static and dynamic analysis to machine learning. The MalOSS framework proposed by Duan et al. (duan2020towards) demonstrates similarities in malicious behaviors across different languages through metadata and API call sequence analysis. Huang et al. utilize graph-based behavior modeling to detect malicious packages (huang2024spiderscan). Zhang et al.(zhang2023malicious) utilize deep learning to understand malicious software’s behavioral semantics, while Liang et al.(liang2023needle) demonstrate the effectiveness of anomaly analysis algorithms. Additionally, GitHub, as the largest source code repository, may serve as a channel for distributing malicious code, with inconsistencies between distributed packages and source code potentially indicating malicious injection (vu2021lastpymile). However, existing detection methods face a critical limitation: detection results are scattered across security blogs and unstructured webpages, failing to reach downstream users effectively. Even with numerous detection tools, malicious packages persist because intelligence remains fragmented. IntelliRadar addresses this through a fundamentally different approach—systematically analyzing malicious package intelligence from dispersed sources, covering both historical and newly reported packages.
@@ 7. Conclusion
In the open-source ecosystem, package managers like PyPI and NPM lack up-to-date intelligence databases, allowing malicious packages to persist. To address this, we developed IntelliRadar, a system leveraging multi-source data collection and LLMs to gather 34,313 pieces of intelligence from 24 sources. The system enables earlier detection of malicious packages, reducing their impact, at \$0.003 per intelligence item. This work presents an efficient solution to enhance open-source software supply chain security.